### v2

# First Measurement of Entanglement Dynamics in the SYK Model using Quantum Computers

Abstract
The Sachdev-Ye-Kitaev (SYK) model, fundamental to quantum chaos, many-body physics, and holographic duality, exhibits both maximal chaos and exact solvability, yet its entanglement entropy has never been measured experimentally. Here, we present the first measurement of entanglement entropy growth under the SYK Hamiltonian using IBM’s superconducting quantum computer as a programmable quantum platform. We implement an optimized swap-based many-body interference protocol tailored for limited qubit connectivity to measure entanglement entropy. Additionally, we enhance the randomized measurement protocol for entropy measurement by expanding the unitary ensemble, which increases the circuit volume and pushes the capability of the near-term devices. However, we tackle this challenge by employing a quantum multi-programming strategy that parallelizes circuit execution. This work paves the way for exploring chaotic quantum dynamics and establishes scalable methods for investigating complex entanglement growth on current quantum platforms.
Introduction.— Quantum entanglement [1] plays a crucial role in various fields, including quantum computing, quantum information science, many-body physics, high-energy and nuclear physics, and black hole physics. Additionally, effectively harnessing quantum information and entanglement dynamics is essential for advancing quantum technologies and exploring complex quantum systems [2, 3]. However, the exponential growth of the Hilbert space associated with increasing system size poses significant challenges for classical computers, making it difficult to compute entanglement entropy for highly entangled states. Besides, experimentally measuring the entanglement entropy and its dynamics in complex quantum systems is challenging due to the specific requirements for state preparation, the execution of time evolution, and the implementation of entropy measurement techniques, such as those based on quantum gas microscopy [4, 5].
One such complex theoretical model for quantum man-body system is the Sachdev-Ye-Kitaev (SYK) model [6, 7, 8, 9], a $0+1$ dimensional quantum mechanical system consisting of $N$ Majorana fermions with random interactions involving $q$ of these fermions at a time, where $q$ is an even number. It is notable for its exact solvability at strong coupling, maximal chaos, and emergent conformal symmetry. Furthermore, the SYK model is a prime example of a strongly interacting quantum non-Fermi liquid lacking a quasiparticle description (for a detailed review of the SYK Model, refer to Ref. [10]). Quantum entanglement plays a crucial role in revealing the nontrivial correlations among the constituents of this representative model [11, 12]. Thus, given the SYK model’s rich physics and the rarity of such solvable, strongly interacting chaotic systems in quantum mechanics, experimentally measuring its quantum entanglement is a valuable endeavor.
In this regard, quantum computers can be utilized as experimental probes to measure the entanglement entropy of the SYK model. In fact, real quantum devices are emerging as highly effective programmable quantum platforms, capable of accurately preparing complex quantum many-body models, and offer a valuable testing ground for novel phenomena that would be difficult or impossible to realize through experimental means [13, 14, 15]. Programmable quantum platforms are being used to study discrete time crystals [16, 17], many-body localization [18, 19, 20, 21], quantum phases, and topologically ordered states [22, 23, 24, 25, 26, 27, 28] in quantum many-body systems. Building on our previous work that investigated the Page Curve entanglement dynamics using quantum computers as programmable platform [29], we now shift our focus to a more complex measurement of entanglement dynamics in the SYK model. Although some studies [30, 31, 32, 33, 34, 35, 36, 37, 38, 39] have addressed the SYK model, we are the first to measure the entanglement entropy within this model using a quantum computer. Specifically, we report our findings on the growth of entanglement entropy for subsystems of varying sizes over time, utilizing IBM’s superconducting quantum computer. This measurement is performed while evolving a total product state with the time evolution operator associated with the SYK Hamiltonian.
However, implementing any quantum protocols on current quantum devices, such as IBM’s quantum processors, presents significant challenges due to the errors and noise inherent in these systems. To address this, we apply various quantum error mitigation (QEM) techniques that have been developed and demonstrate their effectiveness in practical applications [40, 41, 42, 43]. As a result, we accurately measure the variation of entanglement entropy over time during the evolution under the SYK Hamiltonian in quantum computers, even for long circuit depths exceeding 600.
The Model.— The SYK model is given by the Hamiltonian,
$H=i^{\frac{q}{2}}\sum_{1\leq i_{1}<i_{2}<\cdots<i_{q}\leq N}J_{i_{1}i_{2}\cdots i_{q}}\,\chi_{i_{1}}\chi_{i_{2}}\cdots\chi_{i_{q}},$ (1)
where $\chi_{i}$ are the Majorana fermions. Each coefficient $J_{i_{1}i_{2}\cdots i_{q}}$ is taken as a real number drawn from a random Gaussian distribution such that $\langle J_{i_{1}i_{2}\cdots i_{q}}\rangle=0$ and $\langle J^{2}_{i_{1}i_{2}\cdots i_{q}}\rangle=\frac{(q-1)!\mathcal{J}^{2}}{N^{q-1}}$. In this study, we restrict ourselves to $q=4$ as this case represents the dominant low-energy model with time-reversal symmetric interactions.
To implement the SYK Hamiltonian in quantum computers, we use the Jordan-Wigner transformation [44] to map the Majorana fermions into Pauli operators in the following way,
$\chi_{2i-1}\rightarrow Z_{1}Z_{2}\cdots Z_{i-1}X_{i},\,\,\,\chi_{2i}\rightarrow Z_{1}Z_{2}\cdots Z_{i-1}Y_{i}.$ (2)
The Hamiltonian now becomes,
$H=\sum_{n=1}^{M}a_{n}\hat{P}_{n},$ (3)
where $M$ is the number of terms in Eq. (1), $\hat{P}_{n}$ is the Pauli operator of weight $N/2$, i.e. it acts on the Hilbert space of $N/2$ qubits, and $a_{n}$ is the corresponding coefficient.
Here, we introduce the entanglement entropy measure we use in this study. Consider a system of $N$ qubits in a pure state $|\psi\rangle$, with its density matrix given by $\rho=|\psi\rangle\langle\psi|$. If we partition this system into two subsystems consisting of $L$ and $M=N-L$ qubits, the entanglement entropy, specifically the $n$-Rényi entropy associated with the $L$-qubit subsystem, is defined as
$S^{(n)}_{L}=\frac{1}{1-n}\log\left[\mathrm{Tr}(\rho_{L}^{n})\right],$ (4)
where $\rho_{L}$ is the reduced density matrix for the subsystem with $L$ qubits. This reduced density matrix is obtained by tracing out the subsystem of $M=N-L$ qubits, such that $\rho_{L}=\mathrm{Tr}_{M}\rho$. Our focus in this study is $n=2$, corresponding to the 2-Rényi entropy (commonly referred to as Rényi entropy), $S^{(2)}$. Additionally, as $n$ approaches 1, the $n$-Rényi entropy converges to the von Neumann entropy:
$S^{vN}_{L}=-\mathrm{Tr}\left(\rho_{L}\log\rho_{L}\right).$ (5)
To execute the time evolution driven by the SYK Hamiltonian in quantum computers, we use first-order Trotterization [45, 46, 47],
$U(t)=e^{-iHt}\approx\left(\prod_{n=1}^{M}e^{-iH_{n}t/r}\right)^{r},$ (6)
where $H_{n}=a_{n}\hat{P}_{n}$ and $r$ is the number of Trotter steps. The Trotter timestep size is defined as $\delta t=t/r$. We want to emphasize that the SYK Hamiltonian includes all-to-all quartic interactions, which causes the number of terms in the Hamiltonian to grow approximately as $O(N^{4})$ with respect to the number of fermions, $N$. Furthermore, the circuit depth associated with a single Trotter step in first-order Trotterization increases polynomially with $N$. For example, for $N=6$, 8, and 10, the circuit depth $d_{\mathrm{circuit}}$ for a single Trotter step is $d_{\mathrm{circuit}}=84$, 494, and 1726, respectively, before transpilation. Additionally, the gate complexity associated with first-order Trotterization, as shown in Refs. [30], roughly scales as $O(N^{10})$, unless non-trivial optimizations are implemented (for recent works on higher-order Trotterization, see Refs [48, 49]). Consequently, for current quantum computers, where Trotterization is the most appropriate technique for Hamiltonian simulation, the time evolution driven by the SYK Hamiltonian for larger fermion numbers $N>6$ (or $N_{\mathrm{qubit}}>3$) becomes highly complex.
Entanglement Entropy Measuring Protocols.— We consider two protocols to measure the purity and Rényi entropy of subsystems in a total quantum state: 1. swap-based many-body interference (SWAP-MBI) protocol and 2. randomized measurement (RM) protocol.
The swap-based many-body interference protocol [50, 51, 52] is based on the swap test which is presented in the quantum circuit Fig. 1,
{quantikz}
[…]
Figure 1: The quantum circuit associated with the swap-based many-body interference (SWAP-MBI) protocol. Here the Rényi entropy of $L$-qubit subsystem is determined by executing the swap test.
where the purity associated with the $L$-qubit subsystem $\mathrm{Tr}(\rho_{L}^{2})$ can be estimated by $\mathrm{Tr}(\rho_{L}^{2})=2P_{0}-1$ where $P_{0}$ is the probability of finding the ancilla qubit in $|0\rangle$ state. Now, the Rényi entropy $S^{(2)}$ can be readily obtained by $S^{(2)}=-\mathrm{log}(2P_{0}-1)$.
On the other hand, the randomized measurement protocol (RM protocol) [53, 54, 55, 56, 57, 58, 59] relies on the principle that the Rényi entropy of a quantum system is reflected in the statistical correlations between the outcomes of measurements conducted on random bases. After evolving the initial state of $N_{\mathrm{qubit}}$ system with the time evolution operator $U(t)$ to obtain the quantum state $|\psi\rangle$ or the associated density matrix $\rho=|\psi\rangle\langle\psi|$, we focus on measuring the purity and Rényi entropy associated with the subsystem $A$ of $L$ qubits. Now, one applies a product of single-qubit unitary operators
$\hat{U}_{a}=U^{(2)}_{1}\otimes U^{(2)}_{2}\otimes...\otimes U^{(2)}_{L},$ (7)
where each of $U^{(2)}_{i}$ is drawn independently from the circular unitary ensemble (CUE) of the $SU(2)$. Afterward, the measurements of the qubits are done on the computational basis ($Z$-basis). The corresponding quantum circuit is shown in Fig. 2.
{quantikz}
[…]
Figure 2: The quantum circuit associated with the randomized measurement (RM) protocol. Here the local Haar-random unitaries $\hat{U}_{a}$ are applied at $L$-qubit subsystem whose Rényi entropy is being measured.
For each $\hat{U}_{a}$, one conducts repeated measurements to obtain the statistics for estimating the occupation probabilities, $P_{\hat{U}_{a}}(\mathbf{s}_{L})=\mathrm{Tr}[\hat{U}_{a}\rho\hat{U}^{\dagger}_{a}|\mathbf{s}_{L}\rangle\langle\mathbf{s}_{L}|]$ of computational basis states $|\mathbf{s}_{L}\rangle=|s_{1},s_{2},...,s_{L}\rangle$ with $s_{i}=0,1$. Here, $\hat{U}_{a}$ acts only on the subspace of $L$ qubits. Afterward, the entire process is repeated for $N_{U}$ different randomly drawn instances of $\hat{U}_{a}$.
After determining the set of outcome probabilities $P_{\hat{U}_{a}}(\mathbf{s}_{L})$ of the computational basis states $|\mathbf{s}_{L}\rangle$ for one instance of random unitaries, $\hat{U}_{a}$, one computes the following quantity,
$X_{a}=2^{L}\sum_{\mathbf{s}_{L},\mathbf{s}^{\prime}_{L}}(-2)^{-D[\mathbf{s}_{L},\mathbf{s}^{\prime}_{L}]}P_{\hat{U}_{a}}(\mathbf{s}_{L})P_{\hat{U}_{a}}(\mathbf{s}^{\prime}_{L})$ (8)
where $D[\mathbf{s}_{L},\mathbf{s}^{\prime}_{L}]$ is the Hamming distance between bitstrings $\mathbf{s}_{L}=s_{1}s_{2}...s_{L}$ and $\mathbf{s}^{\prime}_{L}=s^{\prime}_{1}s^{\prime}_{2}...s^{\prime}_{L}$, measuring how different two bitstrings are, i. e., $D[\mathbf{s}_{L},\mathbf{s}^{\prime}_{L}]\equiv\#\{i\in A|s_{i}\neq s^{\prime}_{i}\}$. Consequently the ensemble average of $X_{a}$, denoted by $\overline{X}$,
$\overline{X}=\frac{1}{N_{U}}\sum_{a=1}^{N_{U}}X_{a}$ (9)
is the second-order cross-correlations across the ensemble of discrete $N_{U}$ random unitaries $\hat{U}_{a}$, and provides the estimation of the purity $\mathrm{Tr}(\rho_{L}^{2})$ associated with a subsystem of $L$ qubits. Finally, the Rényi entropy is
$S^{(2)}=-\mathrm{log}\overline{X}$ (10)
IBMQ Results.— Despite the importance of the SYK model and the entanglement entropy measurement, implementing them on near-term quantum computers is challenging due to the complexity of the Pauli strings and the computation overhead of the two entanglement entropy measurement protocols, the SWAP-MBI and the RM protocols. In particular, one Trotter step has 126 circuit depths and 39 two-qubit gates (CX or CZ). Since we compute up to five Trotter steps, the largest circuit depth and the number of two-qubit gates are 630 and 195, respectively. Based on the Trotter step implementation, the SWAP-MBI protocol needs 196 additional circuit depth to implement the SWAP test in the protocol. On the other hand, the RM protocol needs only 5 additional circuit depths. However, the protocol needs a large number of repetitions for the random sampling of the $SU(2)$ circuits. To overcome those challenges, we apply several methods detailed below.
We conduct our experiment on the IBM Quantum computer, ibm_marrakesh which has 156 qubits with basis gates CZ, ID, RX, RZ, RZZ, SX, X. We evolve a product state of $N_{\mathrm{qubit}}=3$ that corresponds to $N=6$ Majorana fermions, with time-evolution operator driven by the SYK Hamiltonian via the first-order Trotterization up to $t=10$ (in unit of $\hbar/E$ where $E$ is the overall energy scale of the SYK Hamiltonian) in five Trotter steps (Trotter step size 2.0). In each Trotter step, we employ the entanglement entropy measuring protocols: SWAP-MBI and RM protocols to measure the entropy of left single-qubit and two-qubit subsystems, respectively. As going beyond the case $N>6$ for the SYK model is highly challenging on near-term quantum computers due to the polynomial growth of gate complexity with $N$, we focus on the feasible case of $N=6$ Majorana fermions or $N_{\mathrm{qubit}}=3$ qubits. Additionally, we employed an optimized version of the SWAP-MBI protocol, designed specifically for quantum devices with limited qubit connectivity [29] as shown in Fig. 3.
{quantikz}
[…]
Figure 3: The optimized SWAP-MBI protocol devised for IBM quantum devices with limited qubit connectivity. Here, the quantum circuit for determining the Rényi entropy for the $L=2$ left-subsystem, with $M=N_{\mathrm{qubit}}-2$, is shown as an example.
Still, running quantum algorithms on near-term quantum devices, such as IBM Quantum processors, poses a significant challenge due to the errors and noise present in these systems. But, we precisely measure the variation of entanglement entropy with time under the SYK Hamiltonian evolution in quantum computers up to long circuit depths (more than 600 depths, five timesteps) by applying four quantum error mitigation methods to mitigate the errors and noise inherent to quantum devices: Zero-Noise Extrapolation (ZNE), Pauli Twirling (PT), Dynamic Decoupling (DD), and Matrix-free Measurement Mitigation (M3). Refer to Ref. [43] for further details.
Figure 4: The Rényi entropy $S^{(2)}$ with respect to time $t$ (in arbitrary unit) for the left-subsystem of $L=1$ and $L=2$ qubits (two and four Majorana fermions, respectively), determined using swap-based many-body interference protocol (SWAP-MBI). Here, we set the Trotter stepsize as $\delta t=2$.
During our measurement of the subsystem entanglement entropy over time, we discovered that the SWAP-MBI protocol is more accurate for measuring entropy at early times when it is relatively small. In contrast, the RM protocol, which relies on local Haar-random unitaries, tends to produce incorrect results due to significant statistical fluctuations in its purity estimation during these early times.
As illustrated in Fig. 4, there is excellent agreement between the exact and measured Rényi entropy $S^{(2)}$ on the ibm__\__marrakesh device. The slight deviation observed at larger times (specifically, $t=6-10$) can be attributed to inherent Trotter error. However, at these times, the noiseless Qiskit simulation and the error-mitigated measured values of the SWAP-MBI protocol align very well.
Figure 5: The Rényi entropy $S^{(2)}$ with respect to time $t$ determined using the randomized measurement protocol.
We measure the Rényi entropy using the RM protocol for $t=6,8,$ and $10$, as shown in Fig. 5. Again, we see excellent agreement between exact results and results obtained by RM protocol at ibm__\__marrakesh. A key challenge of the RM protocol is its reliance on an ensemble of local Haar-random unitaries to minimize the variance in entropy computation. In our experiments, we utilize $150$ samples of the RM protocol circuit. Besides, as we apply the Zero Noise Extrapolation (ZNE) and the Pauli Twirling methods to mitigate errors, each method replicates the variation three times (ZNE) and ten times (PT), respectively. Therefore, the total number of circuits for the RM protocol at each measurement time is 4,500 ($150\times 3\times 10$). Since we have three measurement points (6, 8, and 10), the total number of circuits is 135,000.
Therefore, to efficiently utilize the IBMQ computer, we parallelize the execution of quantum circuits through Quantum Multi-Programming (QMP), which allows for the concurrent execution of multiple quantum circuits, regardless of their types or depths of the circuits. The efficiency of the QMP has been demonstrated in various quantum algorithms, such as Grover’s search, quantum amplitude estimation, and quantum support vector machine [60, 61, 62]. Despite the efficiency of the QMP, it does introduce some challenges, such as crosstalk between different circuits. We mitigate these side effects following the technique detailed in Ref. [60]. In our QMP implementation, we parallelized five circuits within a QMP package with one physical idle qubit between circuits. Therefore, we executed 27,000 circuits by parallelizing them by 5 circuits to execute the 135,000 circuits.
Conclusion and Outlook.— Using IBM’s superconducting quantum computer, we have achieved the first experimental measurement of entanglement entropy growth in the SYK model, which marks a significant milestone in the quantum simulation of chaotic many-body systems. Despite the challenges posed by the dense interaction structure of the SYK Hamiltonian and the computational demands of entanglement entropy protocols, we successfully implemented both the SWAP-based many-body interference (SWAP-MBI) and randomized measurement (RM) protocols on IBM’s superconducting quantum hardware. To address the limitations of qubit connectivity, we employed an optimized version of the SWAP-MBI protocol and extracted meaningful entanglement dynamics from quantum circuits of large circuit depths. Importantly, we tackled the substantial resource demands of the RM protocol by utilizing Quantum Multi-Programming (QMP) to parallelize 135,000 circuits. This approach significantly enhanced the efficiency of entropy estimation without compromising accuracy. Our results show excellent agreement with noiseless simulations, validating the SWAP-MBI and RM protocols across different entanglement regimes and timescales.
This work provides a blueprint for measuring entanglement entropy in strongly interacting models on near-term devices and demonstrates how tailored algorithmic and architectural innovations can reveal complex quantum dynamics within the constraints of current quantum hardware. Our approach lays the groundwork for exploring information scrambling, quantum chaos, and holographic dualities in increasingly complex quantum systems.
Acknowledgments
T.A.C would like to thank the High Energy Theory group in the Department of Physics and Astronomy at the University of Kansas for their hospitality and support where this work has been done. R.S.S. is supported by Laboratory Directed Research and Development (LDRD No. 23-051) of Brookhaven National Laboratory and RIKEN-BNL Research Center. This work is supported by the Brookhaven National Laboratory LDRD No. 24-061 (K.Y.). This research used quantum computing resources of the Oak Ridge Leadership Computing Facility, which is a DOE Office of Science User Facility supported under Contract DE-AC05-00OR22725. This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC award DDR-ERCAP0028999 and DDR-ERCAP0033558.